OpenStudy (mimi_x3):

Just for fun. :) \[\int_{0}^{\pi/2}xcot(x)dx\]

hartnn (hartnn):

is the answer \(\pi^3\)/24 ???

hartnn (hartnn):

square root of that

OpenStudy (mimi_x3):

Nope!!!

OpenStudy (mimi_x3):

It cannot be that easy!!

OpenStudy (hba):

(1/2) pi ln(2)

hartnn (hartnn):

oh, yes... \(I=\int_{0}^{\pi/2}x\cot(x)dx \\I=\int_{0}^{\pi/2}(\pi/2-x)\tan(x)dx\) now multiply both equations to eliminate trignometric function

OpenStudy (hba):

@Mimi_x3 Please respond.

hartnn (hartnn):

hba, your answer is correct.

hartnn (hartnn):

did u integrate by parts

OpenStudy (mimi_x3):

Soryy!! Faster than expected!! I think @hba cheated :)

hartnn (hartnn):

you think ? i know :P

OpenStudy (hba):

I did not solve it guys.

OpenStudy (hba):

Some one else solved it for me.

OpenStudy (hba):

:P

OpenStudy (mimi_x3):

It was obv :P impossible to solve it in a couple of seconds.

hartnn (hartnn):

oh, the catch was to add, not multiply :P hba, wolf.

OpenStudy (hba):

Yes @hartnn Yes :P

OpenStudy (hba):

@Callisto Why did you not comment ?

OpenStudy (callisto):

My comment... @hartnn @Mimi_x3 ♥♥♥